# Thermal Management Design and Key Technology Validation for PandaX Underground Experiment

### v1

@@ 4.3 Typical Thermal Load and Access Strategy Analysis
The ACR is relatively inexpensive and has a large cooling power, which can meet most of the cooling requirements of xenon detectors. We will analyze four typical thermal loads, and taking the PandaX-4T detector as an example, provide a simplified thermal management solution to illustrate our design approach and provide reference for future system design.
@@ 4.3.1 Thermal Load with Precision Temperature Control
The cooling power requirements are not large when the PandaX detector and the distillation tower are in 178 K steady-state operation. However, the temperature stability requirement is strict, with a tolerance of ±0.1 K or even ±0.01 K. The “thermal load A” in figure 7 can be directly connected to the centralized cooling system. The design of the precision temperature control cold head based on the ACR is described in section 5, where the low-temperature coolant is the primary cooling source.
@@ 4.3.2 High power thermal load with high temperature
Typical thermal loads include the initial cooling of a large-scale xenon gas storage system and the PandaX detector, which generally start cooling from room temperature to the target temperature over several days. Large cooling power is required regardless of temperature stability. All the ACRs should be in operation. The temperature stability of the coolant needs to be ensured when the other thermal loads with precision temperature control requirements operate in the CCS. A mixer vessel between such thermal loads and the cooling pipeline is useful for these thermal loads, as shown in figure 7.
The mixer is a transition between thermal loads B and the CCS. Functionally, the mixer is similar to the coolant storage tank on a refrigerator. The mixer has four ports for coolant input and output. The coolant driven by the gear pump flows out from port MV-4 at a relatively high flow rate to cool the “thermal load B” and then returns to the mixer through port MV-3. The low-temperature coolant enters the mixer through valve VD and port MV-1 and then returns to the CCS main coolant pipeline through valve VC and port MV-2. Ports MV-1 and MV-3 are close to each other and aligned to achieve a better mixing effect to improve temperature uniformity. The temperature signal is input to the temperature controller, which calculates the opening ratio of the control valve VD based on the upper computer command and the control law to control the coolant in the mixer within the setting temperature range. Though the temperature of the coolant returning to the pipeline through valve VC might be high, the temperature stability of the cooling supply system can still be ensured when the returning flow rate is low. For example, the Rn adsorption unit requires a cooling source of 180$\sim$200 K [15], which can be connected to the CCS through the mixer.
@@ 4.3.4 Thermal Load of Pressure Precision Control of the Detector
During the operation of a dual-phase detector, the significant impact of thermal load is the pressure increase in the detector. The current PandaX cooling systems transfer cooling power from the cold head to the gas xenon through copper cold fingers, and xenon gas is liquefied to reduce the detector pressure. The temperature of the cold head is strictly controlled to ensure the temperature stability in the process [10]. Furthermore, the pressure of the detector can be adjusted by changing the gas phase volume and the total amount of xenon [26]. Therefore, a gas pressure control heat exchanger can be set up based on the convenient supply of coolant, as shown in figure 8, the coolant is at the cold side while the hot side is connected to the gas in the detector via pipeline. It acts as a low-temperature adsorption pump functionally, and the gaseous xenon is absorbed and liquefied, then accumulated to the bottom for gravity and transported to the detector due to gravity or a gear pump. Considering the temperature gradient of heat transfer, the coolant of 163K ensures the equilibrium pressure of 0.1 MPa on the hot side of the heat exchanger. Compared to the rated operation pressure of 0.2 MPa in the detector, the requirement of the flow resistance of the gas pipeline is not strict for the pressure difference of 0.1 MPa. The flow control valve on the gas pipeline can precisely control the gas flow from the detector to the pressure-controlled heat exchanger, thereby obtaining stable detector pressure. For example, the steady-state operation of PandaX-4T requires a cooling power of about 100 W, while the xenon vaporization heat is about 92 J/g, the pressure in the detector can be balanced by liquefying 1 g/s xenon gas in the heat exchanger. This method is the direct measurement and control of the detector pressure [26], which can control the detector pressure more quickly and accurately and reduce the technical requirements of the temperature control compared to the precise temperature control method using the cold head [10]. The gas phase of the detector can also be connected to the inlet of the gas circulation pump via the gas pipeline and flow control valve, realizing the precise pressure control function of the detector. To ensure the liquid xenon temperature entering the detector is constant, a packing structure similar to a distillation column is added to the return liquid xenon pipeline. The packings are located in the gas phase of the detector, and the liquid xenon with lower temperature exchanges heat and mass with the gas xenon on the surface of the packings; some gas xenon is liquefied and quickly balanced to the equilibrium temperature, then injected into the specified position of the detector by gravity, which can keep the stability of the detector operation without the requirement of the complex temperature control system. The packing acts as an adsorption pump functionally for gaseous xenon.
@@ 5 Design of Cold Head based on ACR (CHACR)
The ACR can achieve a temperature of below 160K. However, its operating principles determine the temperature fluctuation of the output is relatively large at the level of ±2 K, which can not meet the temperature stability requirement of 178±0.1 K for the PandaX experiment system. It is feasible to design and manufacture CHACR with reference to PTR and GM cold heads. The refrigerant flow control based on a flow regulating valve is used as the coarse adjustment, precise adjustment is carried out by electric heating, and the control signal is all from the temperature controller. figure 9 shows the design principle of CHACR. The coolant channels are designed at the top of the copper rod with a diameter of 150mm, and a precision temperature sensor is at the bottom of the copper rod for temperature control, and the temperature is defined as the output temperature of CHACR. The bottom of the copper rod is thermally connected to the cold finger via the thermal adhesive, and the cold finger can cool down the xenon gas in the PandaX detector or the distillation tower directly. For stable thermal load, increasing the cold coolant flow rate can decrease the output temperature of the CHACR and vice versa. As a thermally conductive structure, the copper rod itself has a low-pass filtering effect on temperature fluctuations at the input end, so a sampling frequency of 10Hz is sufficient for the CHACR temperature controller. The coolant flow rate should be operated within 10% to 30% of the full precision electric heating power of the temperature controller, which can balance system efficiency and fast response to transient thermal loads. The cooling power of the AL600 cold head with a 127 mm diameter is about 950 W at 178 K. The CHACR can obtain the same cooling power or more.

### v2

@@ 4.3 Typical Thermal Load and Access Strategy Analysis
The ACR is relatively cost-effective and offer substantial cooling power, which can meet most of the cooling requirements of xenon detectors. Four typical thermal loads were analyzed, using PandaX-4T detector as case study, provide a simplified thermal management solution to illustrate design approach and provide a reference for future system designs.
@@ 4.3.1 Thermal Load with Precision Temperature Control
The cooling power requirements are not large when the PandaX detector and the distillation tower are in 178 K steady-state operation. However, the temperature stability requirement is strict, with ±0.1 K or even ±0.01 K tolerance. The “thermal load A” in figure 8 can be directly connected to the CCS. The design of the precision temperature control cold head based on the ACR is described in section 5, where the low-temperature coolant is the primary cooling source.
@@ 4.3.2 High power thermal load with high temperature
Typical thermal loads include the initial cooling of a large-scale xenon gas storage system and the PandaX detector, which generally which typically requires several days to transition from room temperature to the target temperature. A large cooling power is required, regardless of temperature stability. All the ACRs should be in operation. Maintaining temperature stability in the coolant is critical when other thermal loads with precise temperature control requirements are active within the CCS. A mixer vessel connecting these thermal loads to the cooling pipeline is beneficial, as illustrated in figure 8.
The mixer is a transition between thermal loads B and the CCS. Functionally, the mixer is similar to the coolant storage tank on a refrigerator. The mixer has four ports for coolant input and output. The coolant driven by a gear pump flows out from port MV-4 at a relatively high flow rate to cool the “thermal load B” and then returns to the mixer through port MV-3. The low-temperature coolant enters the mixer through valve VD and port MV-1 and then returns to the CCS main coolant pipeline through valve VC and port MV-2. Ports MV-1 and MV-3 positioned closely and aligned to achieve a better mixing effect to improve temperature uniformity. The temperature signal is input to the temperature controller, which calculates the opening ratio of the control valve VD based on the upper computer command and the established control law, ensuring that the coolant in the mixer remains within the designated temperature range. Although the temperature of the coolant returning to the pipeline through valve VC may be high, the temperature stability of the cooling supply system can still be maintained when the returning flow rate is low. For instance, the Rn adsorption unit requires a cooling source of 180$\sim$200 K [15], which can be integrated with the CCS through the mixer.
@@ 4.3.4 Thermal Load of Pressure Precision Control of the Detector
During the operation of a dual-phase detector, the significant impact of thermal load is the pressure increase in the detector. The current cooling systems used in PandaX transfer cooling power from the cold head to the gas xenon through copper cold fingers, enabling the liquification of xenon gas to reduce the detector pressure. The temperature of the cold head is strictly controlled to ensure temperature stability throughout the process [10]. Furthermore, the pressure within the detector can be adjusted by changing the gas phase volume and the total amount of xenon present [26]. To facilitate this, a gas pressure control heat exchanger can be set up based on the convenient supply of coolant, as illustrated in figure 9, the coolant is at the cold side, while the hot side connects to the gas within the detector via pipeline. This setup functions similarly to a low-temperature adsorption pump functionally, and the gaseous xenon is absorbed and liquefied, then accumulated to the bottom for gravity and transported to the detector due to gravity or a gear pump. Considering the temperature gradient during heat transfer, the coolant of 163 K ensures an equilibrium pressure of 0.1 MPa on the hot side of the heat exchanger. When compared to the rated operating pressure of 0.2 MPa in the detector, the requirement of the flow resistance in the gas pipeline is not strict for the pressure difference of 0.1 MPa. The flow control valve on the gas pipeline plays a crucial role in precisely managing the gas flow from the detector to the pressure-controlled heat exchanger, ensuring stable pressure within the detector. For instance, the steady-state operation of PandaX-4T requires a cooling power of about 100 W and the xenon vaporization heat is about 92 J/g. The pressure in the detector can be balanced by liquefying 1 g/s xenon gas in the heat exchanger. This method is the direct measurement and control of the detector pressure [26], which can control the detector pressure more quickly and accurately and reduce the technical requirements of the temperature control compared to the precise temperature control method using the cold head [10]. The gas phase of the detector can also be connected to the inlet of the gas circulation pump via the gas pipeline and flow control valve, thereby enabling precise pressure control function of the detector. To maintain the liquid xenon temperature entering the detector constant, a packing structure similar to a distillation column is added to the return liquid xenon pipeline. The packing is located in the gas phase of the detector, and the liquid xenon with lower temperature exchanges heat and mass with the gas xenon on the surface of the packing; some gas xenon is liquefied and quickly balanced to the equilibrium temperature, then injected into the specified position of the detector by gravity, which can keep the stability of the detector operation without the requirement of the complex temperature control system. The packing acts as an adsorption pump for the gaseous xenon.
@@ 5 Design of Cold Head based on ACR (CHACR)
The ACR can achieve a temperature below 160 K. However, its operating principles determine the temperature fluctuation of the output is relatively large at the level of ±2 K, which cannot meet the temperature stability requirement of 178±0.1 K for the PandaX experiment system. It is feasible to design and manufacture CHACR concerning PTR and GM cold heads. The coolant flow control based on a flow regulating valve is used as the coarse adjustment, precise adjustment is carried out by electric heating, and the control signal is all from the temperature controller. Figure 10 shows the design principle of CHACR.
The coolant channels are designed at the top of the copper rod with a diameter of 150 mm, a precision temperature sensor is at the bottom of the copper rod for temperature control, and the temperature is defined as the output temperature of CHACR. The bottom of the copper rod is thermally connected to the cold finger via the thermal adhesive, and the cold finger can cool down the xenon gas in the PandaX detector or the distillation tower directly. For stable thermal load, increasing the cold coolant flow rate can decrease the output temperature of the CHACR and vice versa. As a thermally conductive structure, the copper rod has a low-pass filtering effect on temperature fluctuations at the input end. Therefore a 10 Hz sampling frequency is sufficient for the CHACR temperature controller. The coolant flow rate should be operated within 10% to 30% of the full precision electric heating power of the temperature controller, which can balance system efficiency and fast response to transient thermal loads. The cooling power of the AL600 cold head with a 127 mm diameter is about 950 W at 178 K. The CHACR can obtain the same cooling power or more.